OpenStudy (anonymous):

\[0.04(x+50)^{2}-3(x+50)+56\] Help me with this problem

OpenStudy (anonymous):

This can be expanded to: 0.04(x^2+100x+2500)-(3x+150)+56 If we expand even further, we get: 0.04x^2+4x+100-3x-150+56 =0.04x^2+x+6 =x^2/25+x+6 The we can factor this into ((x+10)(x+15))/25=0 Then multiply both sides by 25. Now we have (x+10)(x+15)=25 Solving for x, we have; x=-15 or x = -10

OpenStudy (anonymous):

doesn't look like an equation so not suer why you are solving for x

OpenStudy (anonymous):

was this set = 0?

OpenStudy (anonymous):

u

OpenStudy (anonymous):

i mean u=

OpenStudy (anonymous):

the original problem was P(x) and subtract 50 from both sides so its 0 i guess

OpenStudy (anonymous):

was the problem to solve an equation or just to expand? there is not number on the other side of the equal sign. in fact there is no equal sign

OpenStudy (anonymous):

you lost me but that is ok, i will not worry about it